# DistALANER: Distantly Supervised Active Learning Augmented Named Entity Recognition in the Open Source Software Ecosystem

### 25 Feb 2024 v1

@@ 6 Experimental setup
Training and test data preparation: In order to train the NER model in Stage 3, we need training data. For this purpose we select all the bug descriptions from the years 2004 through 2013 inspired by  Hazra et al. (2021b, 2023). These descriptions, nearly 65K in number are automatically annotated using the first two stages of our framework. Thus, the training data only has auto-curated (aka silver) entity labels without any human involvement. For the test set we consider the data from the years 2016 to 2019. To assess the performance of the model, which has been trained using distantly labeled data, we perform human annotation of a subset of the test set.
Human annotation: We present 500 bug descriptions to four domain experts, each claiming over three years of experience in the Ubuntu ecosystem and package management. The resulting inter-annotator agreement among these annotators is 0.625 (see Appendix E for more information). For the Launchpad QA dataset, we employ the same group of experts to annotate the entities in 500 question-answer pairs, ensuring consistency in our data annotation process across both datasets.
@@
Table 3: Recall rate for named entity recognition approaches. The first sub-row of each row shows the results for the Ubuntu bug dataset while the second sub-row shows the results for the Launchpad QA dataset (inference phase only). The best and the second best results in the first sub-row (bug dataset) are highlighted in dark and light green respectively. The best and the second best results in the second sub-row (QA dataset) are highlighted in dark and light purple respectively. BERT-CRF is significantly different from domain specific SoftNER (*p = 0.039)
Metric for evaluation: To evaluate all the models, we compute the recall rate inspired by Tu and Lignos (2021). Recall rate is the proportion of actual entity types in ground truth that are accurately predicted. We use this metric rather than the F1 score in cognizance of the fact that human annotations are sometimes incomplete while the model is able to generate the correct entity type (see Appendix F).
@@ 7 Results
We evaluate all the baselines for both the Ubuntu bug dataset and the Launchpad QA dataset. The main results are noted in Table 3. Our experiments involve two different setups – (a) Human-Induced Training (HIn), and (b) Human-Only Training (HOn). In the case of a Human-Induced (HIn) configuration, we train the models using all the auto-annotated bug descriptions plus 10% (or 50) human-annotated bug descriptions. These examples of human annotated data are incorporated into the training to provide the models with additional knowledge of gold annotations. For Human-Only training (HOn), only 10% of human annotations are used to train the models (see Appendix G). We do not include any auto-annotated data during HOn configuration training. In both setups, 10% of the human-annotated data is used for training, 20% for validation, and 70% for testing.
@@
Further, we examine the trained models’ performance on the Launchpad dataset. Here, we use the trained models (trained using bug descriptions in both HIn and HOn setups) only for inference purposes, i.e., we do not train the models on the Launchpad dataset. The key idea is to test the performance in a zero-shot transfer setup. Here, BERT-CRF exhibits the best overall performance while BERT-NER and Linear-CRF are in the second best position. In Table 3, we report the best model performance in dark purple while light purple denotes the second best. A detailed analysis of the error cases are presented in Appendix H.
Progressive learning: Progressive learning Chatterjee et al. (2017); Moeed et al. (2020) is a strategy where machine learning models receive training in an incremental manner, with larger chunks of data over time. We conduct our progressive learning experiment by dividing our data into 25%, 50%, and 75% segments among equally distributed entities. Figure 4 presents the distribution of Recall rates across varying training data percentage. We limit this experiment to the top three performing models only, providing a focused perspective on how they react to this progressive learning approach. We observe that BERT-CRF continuously gains with additional data finally achieving the best performance with 100% training data.
Figure 4: Recall rate for varying percentage of training data.
8 Additional Experiments
8.1 Relation extraction
Relation extraction (RE) is the most natural follow-up task of NER. In this case study we attempt to identify the effectiveness of NER in solving the RE task in the context of open source software systems. We identify five broad types of relationships: (a) dependency – a dependency relation between two entity types indicates that one entity depends on or relies on the other entity for its proper functioning or execution (e.g. triplets include (sane-utils, Scanner, dependency), (hplip, HP Printer, dependency), (xserver-xorg-video-intel, Intel Graphics Card, dependency), etc.), (b) conflict – conflict relation refers to a situation where two or more software components, packages, or entities cannot coexist or function harmoniously due to incompatibilities, overlapping functionalities, or conflicting configurations (e.g. triplets include (cups, Printers, conflicts), (Keyboard, Input Method Editor (IME), conflicts), etc.), affected version – these indicate which specific versions of a software are affected by reported issues or bugs (e.g. triplets include (Flatbed Scanner, macOS Mojave, affected version), (Error Code 134, sudo dpkg --configure -a, affected version), etc.), cause and effect – these relations refers to cases where one entity (cause) triggers another entity (effect) (e.g. triplets include (Error Code 502, nginx, cause and effect), (Error Code 401, openssh-server, cause and effect), etc.), and interaction/control – corresponds to relations where an entity exerts dynamic influence on/manipulates another entity (e.g. triplets include (apt, install, interaction/control), (docker-ce, run, interaction/control), etc.).
BERT-CRF 0.46 0.59 0.15 0.49 0.45
BERT-NER 0.44 0.59 0.12 0.45 0.46
RoBERTa-CRF 0.44 0.54 0.14 0.49 0.46
Vanilla BERT 0.42 0.51 0.06 0.45 0.48
Table 4: Relation extraction performance (F1 scores).
9 Conclusion
10 Limitation
11 Ethical consideration
Appendix A Heuristics
Sample Bugs Sample Tag Heuristics Conversion
"Some selected error messages from the time of session login"
Messages - (NN, NNS, IN)
Time - (DT, NN, IN
Error ->O
Package - >O
rely on Xvfb with a less than optimal experience."
Less - (DT, JJR, IN) Package ->O
File - (DT, NN, IN)
File - (NNP, NN, IN)
File - (JJ, NN, IN)
Package ->O
Package ->O
Package ->O
Example Prompt for LLM
Extract and tag entities along with start and end index and return
it in json format from the following paragraph into one of the
following entity type: package, operating system, organization,
command, error, extension, peripheral, software component,
architecture. Paragraph: "…."
Table 6: Sample prompts to generate entities in LLM.
Appendix B Error analysis
NER can help extract entities like library names, function names, software versions, bug IDs, or even contributor names from this sea of information. However, it is vital to ensure the NER system’s accuracy, and this is where error analysis comes in. In this section, we analyze the incorrect predictions and group them into the following types (see Table 8 for examples).
  • Ambiguity errors: These occur when a token could be identified as multiple entity types or when it is unclear whether a token is an entity. For example, ‘Apple’ could refer to the OS or an organization in a bug. Enhancing training data to better cover such ambiguous instances or using more advanced models that are adept at handling ambiguities can help.
  • Boundary detection errors: These occur when the system inaccurately determines the start and end of an entity. For example, the system might identify “Windows NT3.‘..’" as separate entities instead of a single entity referring to a specific Windows version. Solutions can include using techniques like BIO tagging (Begin, Inside, Outside) or increasing the range of training examples with varying entity lengths.
  • Cohesion errors: These are related to the semantics of software domain language. For instance, ‘sudo’, ‘apt’, ‘update’ should be recognized as individual instead of whole ‘sudo apt update’. This problem can be mitigated by employing sequence tagging models and enhancing the training set with more examples of such entities.
  • Errors due to homonyms: Some words may have different meanings depending on the context. For example, ‘Java’ can refer to a programming language, an island, or a type of coffee. Models need to be trained to understand the context to disambiguate the intended meaning.
  • Multi-annotator disagreement: In the domain of open source software, disagreements among multiple expert annotators can frequently arise due to the inherent complexity and variety of software artifacts. These disagreements can be a result of different interpretations of code, differing understanding of software architecture, or varying familiarity with the programming languages and software systems in question. This heterogeneity in expert opinions can pose significant challenges in tasks such as bug labeling, feature identification, or documentation generation.
Appendix C Source details
  • Packages: We use all the packages assembled by the Hazra et al. (2021a) of the paper.
  • Error Codes: We collect these from the Ubuntu page  [^15] https://wiki.ubuntu.com/error%20and%20warning%20messages, and use TagMe (with our active learning based approach) to augment the list with additional error codes.
  • File Extensions: We collect this data from the Wikipedia page  [^16] https://en.wikipedia.org/wiki/Filename_extension.
Appendix D Regex details
Appendix E Interannotator agreement
Appendix F Extended results
In evaluating NER models for lengthy software-related texts, we emphasize the importance of recall for two main reasons:
  • Missed entities carry a high cost: When dealing with software-related texts, overlooking an entity can lead to the omission of vital data. For instance, not recognizing an error code, a software package, or a particular function name can change the outcome from comprehending and addressing a problem to failing to do so. This makes capturing as many relevant entities as we can, which recall assesses, crucial.
    Importance of precise information extraction: Software-related documents typically present crucial details that require precise extraction for subsequent applications like bug tracking, requirement analysis, or code generation. Overlooking an entity in these contexts can be impactful, underscoring the importance of recall as a key metric.
However for complete picture Table 7 presents the precision and macro F1.
@@ Appendix F Extended results
(’apple’, marked as([’O’, ’OS’, ’Organization’])),
(’exit’, marked as ([’Command’, ’O’, ’Software_Component’, ’Package’])
(’ask’, marked as ([’O’, ’Package’]))
@@
For Command - “sudo apt update”
our model marks them separately as “sudo”, “apt” & “update”, but human marked them as a whole.
Table 8: Error analysis examples.
Appendix G Motivation of different split of HOn
Method ARC CMD ERR EXT OS ORG PKG PRP SOC Overall
Appendix H Hyperparameter settings
@@
Appendix I Issues in finding distant labels using LLMs
In our current research, we employ a consistent instruction-based prompt to extract and tag entities from a given paragraph (see Table 6) for sample prompts). The goal is to return these tagged entities, along with their start and end indices, in a JSON format. The paragraph is analyzed to identify the nine entities. We evaluate the outputs from three different language models: GPT-3.5-Turbo, GPT-4, Google BARD and UniversalNER. Through our analysis, we make several observations: (a) GPT-4 fails to generate the complete JSON format, unlike GPT-3.5-Turbo and Google BARD; (b) The output pattern from GPT-4 and Google BARD exhibits more consistency than that of GPT-3.5-Turbo. For instance, when generating entities using GPT-4 and Google BARD, the model returns the entity, its type, and its start and end indices. However, GPT-3.5-Turbo occasionally returns the entity type and indices without the text itself, or merely refers to specific lines in the text leading to inconsistencies in its output structure; (c) In terms of content, both GPT-3.5-Turbo and GPT-4 extract more accurate entities compared to Google BARD. However, the indices produced by GPT-3.5-Turbo and GPT-4 tend to contain more errors compared to those generated by Google BARD; (d) We notice that all three models sometimes invent their own entity types, which they subsequently return in the output; (e) In the recently published UniversalNER Zhou et al. (2023), the model struggles to generate accurate entities, despite being trained on relevant data. (e) Finally, access to these model predictions are far more expensive compared to our distantly supervised method. See Tables 11 and 12 for some examples.

### 15 Mar 2024 v3

@@ 8 Experimental setup
Training and test data preparation: In order to train the NER model in Stage 3, we need training data. For this purpose we select all the bug descriptions from the years 2004 through 2013 inspired by  Hazra et al. (2021, 2023a, 2023b). These descriptions, nearly 65K in number are automatically annotated using the first two stages of our framework. Thus, the training data only has auto-curated (aka silver) entity labels without any human involvement. For the test set we consider the data from the years 2016 to 2019. To assess the performance of the model, which has been trained using distantly labeled data, we perform human annotation of a subset of the test set.
Human annotation: We present 500 bug descriptions to four domain experts, each claiming over three years of experience in the Ubuntu ecosystem and package management. The resulting inter-annotator agreement among these annotators is 0.625 (see Appendix C for more information). For the Launchpad QA dataset, we employ the same group of experts to annotate the entities in 500 question-answer pairs, ensuring consistency in our data annotation process across both datasets.
@@
Table 4: Recall rate for named entity recognition approaches. The first sub-row of each row shows the results for the Ubuntu bug dataset while the second sub-row shows the results for the Launchpad QA dataset (inference phase only). The best and the second best results in the first sub-row (bug dataset) are highlighted in dark and light green respectively. The best and the second best results in the second sub-row (QA dataset) are highlighted in dark and light purple respectively. BERT-CRF is significantly different from domain specific SoftNER (*$p=0.039$). Table 5 compares methods for HIn based on other metrics.
Metric for evaluation: To evaluate all the models, we compute the recall rate inspired by Tu and Lignos (2021). Recall rate is the proportion of actual entity types in ground truth that are accurately predicted. We use this metric rather than the F1 score in the main table in cognizance of the fact that human annotations are sometimes incomplete while the model is able to generate the correct entity type (see Section 9). Precisely, when dealing with software-related texts, overlooking an entity can lead to the omission of vital data. For instance, not recognizing an error code, a software package, or a particular function name can change the outcome from comprehending and addressing a problem to failing to do so. This makes capturing as many relevant entities as we can, which recall assesses, crucial. Further, software-related documents typically present crucial details that require exhaustive extraction for subsequent applications like bug tracking, requirement analysis, or code generation. Overlooking an entity in these contexts can have an adverse impact, underscoring the importance of recall as a key metric. Nevertheless, in order to make the results complete we also subsequently measure and report the precision, recall and F1 score.
@@ 9 Results
We evaluate all the baselines for both the Ubuntu bug dataset and the Launchpad QA dataset. The main results are noted in Table 4. Our experiments involve two different setups – (a) Human-Induced Training (HIn), and (b) Human-Only Training (HOn). In the case of a Human-Induced (HIn) configuration, we train the models using all the auto-annotated bug descriptions plus 10% (or 50) human-annotated bug descriptions. These examples of human annotated data are incorporated into the training to provide the models with additional knowledge of gold annotations. For Human-Only training (HOn), only 10% of human annotations are used to train the models (see subsection 9.2). We do not include any auto-annotated data during HOn configuration training. In both setups, 10% of the human-annotated data is used for training, 20% for validation, and 70% for testing.
@@
Further, we examine the trained models’ performance on the Launchpad dataset. Here, we use the trained models (trained using bug descriptions in both HIn and HOn setups) only for inference purposes, i.e., we do not train the models on the Launchpad dataset. The key idea is to test the performance in a zero-shot transfer setup. Here, BERT-CRF exhibits the best overall performance while BERT-NER and Linear-CRF are in the second best position. In Table 4, we report the best model performance in dark purple while light purple denotes the second best. A detailed analysis of the error cases are presented in section 12.
@@
9.1 Progressive learning
Progressive learning Chatterjee et al. (2017); Moeed et al. (2020) is a strategy where machine learning models receive training in an incremental manner, with larger chunks of data over time. We conduct our progressive learning experiment by dividing our data into 25%, 50%, and 75% segments among equally distributed entities. Figure 4 presents the distribution of Recall rates across varying training data percentage. We limit this experiment to the top three performing models only, providing a focused perspective on how they react to this progressive learning approach. We observe that BERT-CRF continuously gains with additional data finally achieving the best performance with 100% training data.
Figure 4: Recall rate for varying percentage of training data.
9.2 Motivation of different split of HOn
Method ARC CMD ERR EXT OS ORG PKG PRP SOC Overall
10 Issues in finding distant labels using LLMs
In our current research, we employ a consistent instruction-based prompt to extract and tag entities from a given paragraph (see Table 9) for sample prompts). The goal is to return these tagged entities, along with their start and end indices, in a JSON format. The paragraph is analyzed to identify the nine entities. We evaluate the outputs from three different language models: GPT-3.5-Turbo, GPT-4, Google BARD and UniversalNER. Through our analysis, we make several observations: (a) GPT-4 fails to generate the complete JSON format, unlike GPT-3.5-Turbo and Google BARD; (b) The output pattern from GPT-4 and Google BARD exhibits more consistency than that of GPT-3.5-Turbo. For instance, when generating entities using GPT-4 and Google BARD, the model returns the entity, its type, and its start and end indices. However, GPT-3.5-Turbo occasionally returns the entity type and indices without the text itself, or merely refers to specific lines in the text leading to inconsistencies in its output structure; (c) In terms of content, both GPT-3.5-Turbo and GPT-4 extract more accurate entities compared to Google BARD. However, the indices produced by GPT-3.5-Turbo and GPT-4 tend to contain more errors compared to those generated by Google BARD; (d) We notice that all three models sometimes invent their own entity types, which they subsequently return in the output; (e) In the recently published UniversalNER Zhou et al. (2023), the model struggles to generate accurate entities, despite being trained on relevant data. (e) Finally, access to these model predictions are far more expensive compared to our distantly supervised method. See Tables 11 and 12 (from Appendix) for some examples.
11 Additional experiments for task-based evaluation
11.1 Relation extraction
Relation extraction (RE) is the most natural follow-up task of NER. In this case study we attempt to identify the effectiveness of NER in solving the RE task in the context of open source software systems. We identify five broad types of relationships: (a) dependency – a dependency relation between two entity types indicates that one entity depends on or relies on the other entity for its proper functioning or execution (e.g. triplets include (sane-utils, Scanner, dependency), (hplip, HP Printer, dependency), (xserver-xorg-video-intel, Intel Graphics Card,
dependency), etc.), (b) conflict – conflict relation refers to a situation where two or more software components, packages, or entities cannot coexist or function harmoniously due to incompatibilities, overlapping functionalities, or conflicting configurations (e.g. triplets include (cups, Printers, conflicts), (Keyboard, Input Method Editor (IME), conflicts), etc.), affected version – these indicate which specific versions of a software are affected by reported issues or bugs (e.g. triplets include (Flatbed Scanner, macOS Mojave, affected version), (Error Code 134, sudo dpkg --configure -a, affected version), etc.), cause and effect – these relations refers to cases where one entity (cause) triggers another entity (effect) (e.g. triplets include (Error Code 502, nginx, cause and effect), (Error Code 401, openssh-server, cause and effect), etc.), and interaction/control – corresponds to relations where an entity exerts dynamic influence on/manipulates another entity (e.g. triplets include (apt, install, interaction/control), (docker-ce, run, interaction/control), etc.).
BERT-CRF 0.46 0.59 0.15 0.49 0.45
BERT-NER 0.44 0.59 0.12 0.45 0.46
RoBERTa-CRF 0.44 0.54 0.14 0.49 0.46
Vanilla BERT 0.42 0.51 0.06 0.45 0.48
Table 7: Relation extraction performance (F1 scores).
12 Error analysis
NER can help extract entities like library names, function names, software versions, bug IDs, or even contributor names from this sea of information. However, it’s vital to ensure the NER system’s accuracy, and this is where error analysis comes in. In this section, we analyze the incorrect predictions and group them into the following types.
  • Ambiguity errors: These occur when a token could be identified as multiple entity types or when it’s unclear whether a token is an entity. For example, "Apple" could refer to the OS or an organization in a bug. Enhancing training data to better cover such ambiguous instances or using more advanced models that are adept at handling ambiguities can help.
  • Boundary detection errors: These occur when the system inaccurately determines the start and end of an entity. For example, the system might identify "Windows NT3.‘..’" as separate entities instead of a single entity referring to a specific version of the Windows version. Solutions can include using techniques like BIO tagging (Begin, Inside, Outside) or increasing the range of training examples with varying entity lengths.
  • Cohesion errors: These are related to the semantics of software domain language. For instance, "sudo, apt, update" should be recognized as individual instead of whole "sudo apt update". This problem can be mitigated by employing sequence tagging models and enhancing the training set with more examples of such entities.
  • Errors due to homonyms: Some words may have different meanings depending on the context. For example, "Java" can refer to a programming language, an island, or a type of coffee. Models need to be trained to understand the context to disambiguate the intended meaning.
  • Multi Annotator disagreement: In the domain of open source software, disagreements among multiple expert annotators can frequently arise due to the inherent complexity and variety of software artifacts. These disagreements can be a result of different interpretations of code, differing understanding of software architecture, or varying familiarity with the programming languages and software systems in question. This heterogeneity in expert opinions can pose significant challenges in tasks such as bug labeling, feature identification, or documentation generation.
@@ 12 Error analysis
(’apple’, marked as([’O’, ’B-OS’, ’B-Organization’])),
(’exit’, marked as ([’B-Command’, ’O’, ’B-Software_Component’, ’I-Package’])
(’ask’, marked as ([’O’, ’B-Package’, ’I-Package’]))
@@
For Command - “sudo apt update” our model marks
them separately as “sudo”, “apt” & “update”, but human marked them as a whole.
Table 8: Error analysis.
13 Conclusion
14 Limitation
15 Ethical consideration
Appendix A Prompt
Example Prompt for LLM
Extract and tag entities along with start and end index and return
it in json format from the following paragraph into one of the
following entity type: package, operating system, organization,
command, error, extension, peripheral, software component,
architecture. Paragraph: "…."
Table 9: Sample prompts to generate entities in LLM.
Appendix B Regex details
Appendix C Interannotator agreement
Appendix D Hyperparameter settings